# Improving variational counterdiabatic driving with weighted actions and computer algebra

### v1

@@ Appendix D Details of the computer algebra for general spin-1/2 systems
(6) Compute the commutators $[[H(\lambda)-E_{\lambda}^{(K)}]^{K},A_{\mu}]$ for $\mu=1,\dots,M$ and the double commutators $[[[H(\lambda)-E_{\lambda}^{(K)}]^{K},A_{\mu}],A_{\nu}]$ for all $\mu\leq\nu$.
@@
The total computation time for Algorithm 3 scales as $O(N^{K})$. Concretely, the computation time for steps (1) and (4)–(9) are the same as for the corresponding steps in Algorithm 1. The computation for step (2) takes $O(N^{K-1})$ time, and step (3) takes $O(N^{0})$ time, making them negligible compared to the rest of the algorithm.
Algorithm 4.—This algorithm assumes that the Hamiltonian has the form of Eq. (19). We use symbols and definitions in Appendices B.3 and C.2.
@@
(1) Compute the operators $\tilde{F}_{g}$ for $g=1,2,\dots$ by multiplication and addition of $F_{\gamma}$’s.
(2) Compute the traces $\tilde{\omega}_{g}=\mathop{\mathrm{tr}}(\tilde{F}_{g}*I)$ for $g=1,2,\dots,$ and $\tilde{\omega}_{gg^{\prime}}=\mathop{\mathrm{tr}}(\tilde{F}_{g}*\tilde{F}_{g^{\prime}})$ for $g,g^{\prime}=1,2,\dots$.
(3) For $g=1,2,\dots$, compute the commutators $[\tilde{F}_{g},A_{\mu}]$ for $\mu=1,\dots,M$ and the double commutators $[[\tilde{F}_{g},A_{\mu}],A_{\nu}]$ for all $\mu\leq\nu$.
(4) For all $g,g^{\prime}=1,2,\dots$, compute $\tilde{r}_{\mu,gg^{\prime}}$ by Eq. (69) for all $\mu$, and compute $\tilde{Q}_{\mu\nu,gg^{\prime}}$ by Eq. (70) for all $\mu\leq\nu$.
@@
(5) Compute the scalar coefficient $\tilde{f}_{g}(\lambda)$ for $g=1,2,\dots$ by multiplying $f_{\gamma}(\lambda)$’s, and compute $\partial_{\lambda}\tilde{f}_{g}(\lambda)$ for $g=1,2,\dots$ using the Leibniz rule.
@@
(7) Minimize the function $\Omega_{\lambda}^{(K)}(E)$ to determine $E_{\lambda}^{(K)}$. The details of the minimization are discussed in Appendix C.1.
(8) Compute the scalar coefficients $p_{k}(\lambda)=\binom{K}{k}(E_{\lambda}^{(K)})^{K-k}$ for $k=0,\dots,K$.
(9) Compute $r_{\mu}^{(K)}(\lambda)$ for all $\mu$ and $Q_{\mu\nu}^{(K)}(\lambda)$ for all $\mu\leq\nu$ by Eq. (58).
@@
The computation time for Algorithm 4 scales as $O(N^{K})$, similarly to the above three algorithms.

### v3

@@ Appendix C Details of the computer algebra for general spin-1/2 systems
(6) Compute the commutators $H(\lambda)-E_{\lambda}^{(K)}]^{K},A_{\mu}]$ simultaneously for all $\mu$.
(7) For each $\mu$, compute the commutators $H(\lambda)-E_{\lambda}^{(K)}]^{K},A_{\mu}],A_{\nu}]$ simultaneously for all $\nu=1,\dots,\mu$.
@@
The total computation time for Algorithm 3 scales as $O(N^{K})$ for the algebraic computation (steps (1)–(9)) and $O(M^{3})$ or $O(RM^{2})$ for the scalar computation (step (10)), similarly to Algorithm 1. Concretely, the computation time for steps (1) and (4)–(10) are the same as for the corresponding steps in Algorithm 1. The computation for step (2) takes $O(N^{K-1})$ time, and step (3) takes $O(N^{0})$ time, making them negligible compared to the rest of the algorithm.
Algorithm 4.—This algorithm assumes that the Hamiltonian has the form of Eq. (19). We use symbols and definitions in Appendices A.4 and B.3.
@@
(1) Compute the operators $\tilde{F}_{g}$ for all $g$ by multiplication and addition of $F_{\gamma}$’s.
(2) Compute the traces $\tilde{\omega}_{g}=\mathop{\mathrm{tr}}(\tilde{F}_{g}*I)$ for all $g$ and $\tilde{\omega}_{gg^{\prime}}=\mathop{\mathrm{tr}}(\tilde{F}_{g}*\tilde{F}_{g^{\prime}})$ for all $(g,g^{\prime})$.
(3) For every $g$, compute the commutators $\tilde{F}_{g},A_{\mu}]$ simultaneously for all $\mu$.
(4) For every $g$ and every $\mu$, compute the commutators $\tilde{F}_{g},A_{\mu}],A_{\nu}]$ simultaneously for all $\nu=1,\dots,\mu$.
(5) For all $(g,g^{\prime})$, compute $\tilde{r}_{\mu,gg^{\prime}}$ by Eq. (68) for all $\mu$, and compute $\tilde{Q}_{\mu\nu,gg^{\prime}}$ by Eq. (69) for all $(\mu,\nu)$ such that $\mu\geq\nu$.
@@
(6) Compute the scalar coefficient $\tilde{f}_{g}(\lambda)$ for all $g$ by multiplying $f_{\gamma}(\lambda)$’s, and compute $\partial_{\lambda}\tilde{f}_{g}(\lambda)$ for all $g$ using the Leibniz rule.
@@
(8) Minimize the function $\Omega_{\lambda}^{(K)}(E)$ to determine $E_{\lambda}^{(K)}$. The details of the minimization are discussed in Appendix B.2.
(9) Compute the scalar coefficients $p_{k}(\lambda)=\binom{K}{k}(-E_{\lambda}^{(K)})^{K-k}$ for $k=0,\dots,K$.
(10) Compute $r_{\mu}^{(K)}(\lambda)$ for all $\mu$ and $Q_{\mu\nu}^{(K)}(\lambda)$ for all $\mu\geq\nu$ by Eq. (46).
@@
The computation time of Algorithm 4 scales as $O(N^{K})$ for the first stage and $O(M^{3})$ or $O(RM^{2})$ for the second stage, similarly to the above three algorithms.
Supplemental Material for
Improving variational counterdiabatic driving with weighted actions and computer algebra
Naruo Ohga and Takuya Hatomura
This Supplemental Material contains two figures referenced in the main text.
Figure S1: Driving coefficients $\alpha^{(K)}_{\mu}(\lambda)$ obtained from the weighted variational method for representative systems with $N=12$. The index $\mu$ of driving terms is arranged as shown in the top, with $1\leq\mu\leq 12$ representing one-body terms and $13\leq\mu\leq 29$ representing two-body terms. Each panel shows $\alpha^{(K)}_{\mu}(\lambda)$ for two or three terms $\mu$. Solid curves are for the smallest $\mu$, and dashed curves are for the second smallest $\mu$. Dash-dotted curves are for the third, if it exists. (a) Ferromagnetic system with the one-body driving, for which some of the curves have been shown in Fig. 2(b) in the main text. (b) Antiferromagnetic system with the two-body driving. (c) Spin-glass system with the one-body driving.
Figure S2: Performance of the weighted variational method in representative antiferromagnetic and spin-glass systems of $N=12$. Similar plots for ferromagnetic systems have been shown in Fig. 3 in the main text. Panels (a) and (b) show an antiferromagnetic system with the one-body driving, panels (c) and (d) show the same antiferromagnetic system with the two-body driving, and panels (e) and (f) are for a spin-glass system with the one-body driving. In panels (a), (c), and (e), we plot the time evolution of the fidelity to the ground state $\mathcal{F}^{(K)}(t)$, with the insets showing the close-up of the curves at later times. In panels (b), (d), and (f), we plot the final fidelity $\mathcal{F}_{\mathrm{f}}^{(K)}$ over varied protocol durations $t_{\mathrm{d}}$, where the insets are the zoom-up for short durations. In the antiferromagnetic systems with the one-body driving (panels (a) and (b)), the time evolution without CD driving ($K=\varnothing$) performs better than the time evolution with variational CD driving for small $K$, as discussed in the main text.